OpenStudy (anonymous):

Solve dy/dx=(2cos^2(x)-sin^2(x)+y^2)/2cos(x) given that y=sin(x)+1/v(x) Solve via general substitution

OpenStudy (anonymous):

is this the problem? \[\frac{dy}{dx}=\frac{2\cos^2x-\sin^2x+y^2}{2cosx} \] and \[y=\frac{sinx+1}{v(x)}\]

OpenStudy (anonymous):

yes that is the problem